# Design Principle Transfer in Neural Architecture Search via Large Language Models

### v1

@@ Abstract
Transferable neural architecture search (TNAS) has been introduced to design efficient neural architectures for multiple tasks, to enhance the practical applicability of NAS in real-world scenarios. In TNAS, architectural knowledge accumulated in previous search processes is reused to warm up the architecture search for new tasks. However, existing TNAS methods still search in an extensive search space, necessitating the evaluation of numerous architectures. To overcome this challenge, this work proposes a novel transfer paradigm, i.e., design principle transfer. In this work, the linguistic description of various structural components’ effects on architectural performance is termed design principles. They are learned from established architectures and then can be reused to reduce the search space by discarding unpromising architectures. Searching in the refined search space can boost both the search performance and efficiency for new NAS tasks. To this end, a large language model (LLM)-assisted design principle transfer (LAPT) framework is devised. In LAPT, LLM is applied to automatically reason the design principles from a set of given architectures, and then a principle adaptation method is applied to refine these principles progressively based on the new search results. Experimental results show that LAPT can beat the state-of-the-art TNAS methods on most tasks and achieve comparable performance on others.
@@ Introduction
  • An LLM-assisted framework is proposed to implement the design principle transfer across different NAS tasks, which offers at least three advantages: (i) Learning of the general design principles based on LLM; (ii) Task-specific principle adaptation against domain shit; (iii) Improved interpretability of search space refinement.
@@
To reduce the computational cost, transfer learning has been incorporated into NAS, giving rise to TNAS (Huang et al. 2022; Lee, Hyung, and Hwang 2021; Lu et al. 2020, 2021; Elsken et al. 2020; Zhou et al. 2023). In TNAS, the knowledge gained during the architecture search for one task is applied to aid in the architecture design for others. Depending on the types of transferred knowledge, TNAS can be classified into two categories. One category transfers already found architectures from previous tasks to directly build architectures for new tasks. For instance, Lu et al. (Lu et al. 2020) design architectures for CIFAR-100 and use them to classify images from ImageNet directly. However, the generalization performance of these methods has been criticized. To solve this drawback, multitasking evolutionary NAS (MTNAS) (Zhou et al. 2023) applies these high-performing architectures to guide the architecture search for new tasks instead of solving them directly.
The second category reuses the NAS model built in previous tasks to guide the architecture search for a new task. For example, Lu et al. (Lu et al. 2021) leverage the supernet that has been initialized on ImageNet to improve the search efficiency of evolutionary NAS (ENAS) on new image classification tasks. In another related work (Elsken et al. 2020), a meta-architecture is acquired from previous tasks, and this architecture can be adapted to a new task just through a few gradient-based steps. Huang et al. (2022) fine-tune the performance predictor developed for the Jigsaw task and use it to reduce the search cost on other computer vision tasks. However, these methods still search in a very large search space, consuming a lot of computing resources for obtaining high-quality architectures. In this work, a new paradigm of TNAS is proposed, i.e., design principle transfer. In this method, general design principles are summarized from established architectures and further reused to reduce the search space for new NAS tasks.
@@
An overview of LAPT is presented in Figure 1 and Algorithm 1. This framework consists of two stages, i.e., design principle learning and the architecture search for new tasks. In the design principle learning stage, a set of well-performing architectures $\mathcal{A}$ is collected from a pre-defined search space $\Omega$ (line 1 in Algorithm 1). Then, a prompt is devised to help a pre-trained LLM reason general design principles $P$ from $\mathcal{A}$ (line 2 in Algorithm 1). These learned principles $P$ are transferred as the initial design principles $\{P_{n}\}_{n=1}^{N}$ to help solve the new NAS tasks $\{\mathcal{T}_{n}\}_{n=1}^{N}$ (lines 3-10 in Algorithm 1). Specifically, for the task $\mathcal{T}_{n}$, $P_{n}$ is translated to the subsets $\{\mathcal{C}^{r}_{i},~{}\mathcal{S}^{r}_{i}\}_{i}^{L}$ by the pre-trained LLM, leading to the generation of a refined search space $\Omega_{r}$. Then, a NAS method is applied to search for the promising architectures from $\Omega_{r}$ for $\mathcal{T}_{n}$. These found architectures are applied to adapt $P_{n}$ to $\mathcal{T}_{n}$. These steps related to principle adaptation and architecture search are repeated until reaching a predefined stopping criterion. Finally, the best architecture found in the search process is used to solve $\mathcal{T}_{n}$.
@@
1:  Build an archive $\mathcal{A}$ consisting of performing architectures;
2:  $\{P_{n}\}_{n=1}^{N}\leftarrow$ Prompt the LLM to learn design principles from $\mathcal{A}$;
3:  $Base\leftarrow 0$;
4:  for $n$ from 1 to $N$ do
@@
9:        $P_{n},~{}Base\leftarrow$ Adapt $P_{n}$ based on $\mathcal{B}_{n}$, $Base$, $Best$ and return the updated results; #Alg. 2
@@
Firstly, the pre-trained LLM benefits from exposure to a wide array of programming languages, allowing it to gain awareness of the neural architecture from source codes (Zheng et al. 2023). Nevertheless, due to the token limitation, it becomes infeasible to feed all architecture source codes directly into the LLM. To tackle this issue, Python classes that can instantiate an architecture based on its architectural parameters, i.e., ${\cup}_{i=1}^{L}\{l_{i},~{}s_{i}\}$, are set as prompts. This approach enables LLMs to assimilate knowledge about these neural architectures solely through a few architecture parameters. A simple example of the Python code-based prompt is shown in Figure. 2.
@@
In the architecture search stage, principles $P$ learned in the last stage are used to help build the refined search space $\Omega_{r}$. Specifically, we prompt the pre-trained LLM to translate $P$ into the most available candidate operators $\mathcal{C}_{i}$ and information sources $\mathcal{S}_{i}$ for each layer. In this way, $\Omega_{r}$ can be built following equation (3). Then, the NAS method is used to search architectures from $\Omega_{r}$, aiming to solve the problem shown in equation (1).
However, due to domain shift where the architecture performs differently on different tasks, $\Omega_{r}$ may not be optimal for all the tasks. To alleviate the potential negative effects of domain shift, this work adapts $P$ to the target task based on the newly found architectures. As a result, a task-specific subspace can be built progressively. Specifically, the NAS method is applied to find architecture from $\Omega_{r}$ in an iterative way. In each iteration, if the better-performing architectures are found, LLM is prompted to update $P_{n}$ based on these architectures; otherwise, LLM is required to describe effects of other available candidate operators and information sources that are not in $P_{n}$, promoting the exploration for other promising regions in $\Omega$. Details of this adaption strategy can be seen in Algorithm 2.
@@
Input: design principles $P_{n}$, found architectures $\mathcal{B}_{n}$ and $\mathcal{A}$, $Base$, $Best$, # of selection architectures $r$.
@@
Trans101. Architectures in Trans101 follow the same cell-based structure in NAS201, but there are only 4 candidate operators. Thus, only 4K candidate architectures are included in the space.
@@
Experiments on NAS201. In these experiments, we collect 50 top-performing architectures on ImageNet from NAS201 as the input for LLM. Then, learned principles are used to constrain the range of available operators on each layer, leading to a refined search space of NAS201. We search for individual architecture to solve the image classification task on CIFAR-10 and CIFAR-100 from the refined search space.
@@
In these experiments, GPT-4 is used as the pre-trained LLM for design principle learning and adaptation. Related prompts can be seen in Appendix II. In addition, the effectiveness of different LLMs is shown in Appendix V
Table 2: Comparison with TNAS and NAS methods on NAS201.We present the accuracy achieved on two unseen datasets. Additionally, we provide the number of neural architectures (Archs.) that are trained to achieve accuracy. We run LAPT 20 times with different random seeds and average values are reported.
@@
Comparisons on NAS201. We choose well-known NAS methods and TNAS methods as baselines: (1) the vanilla NAS method REA and the state-of-the-art NAS method HEBO (Cowen-Rivers et al. 2022); (2) the state-of-the-art NAS TNAS methods include TNAS-BO (Shala et al. 2023) and MetaD2A (Lee, Hyung, and Hwang 2021). In LAPT, REA is used for architecture search, and its results (LAPT-REA) are shown in Table 2.
@@
Compared with TNAS methods, benefiting from the search space refinement, LAPT can achieve the SOTA performance. Specifically, all of these methods can reach the top architecture on CIFAR-10, but LAPT is about 25x faster than MetaD2A and 8x faster than TNAS-BO. On CIFAR-100, LAPT outperforms MetaD2A obviously and is about 8 faster than TNAS-BO.
@@
From Table 3, our LAPT can achieve better search performance and search efficiency by a large margin than the NAS methods. Specifically, LAPT can beat BONAS on all the 6 tasks. LAPT outperforms REA on 5 out of 6 tasks and achieve the highest average mode rank.
@@
Table 4: Performance comparison on CIFAR-10 using the search space DARTs. LAPT runs 3 times.
@@
Search results on ImageNet are shown in Table 5. From this table, we can see that because of the search space refinement, LAPT can achieve similar search results with other NAS methods but with about 2$\times$ less search cost (2 GPU days), which is more time-efficient. The learned general principles for DARTs are shown in Appendix IV.
@@
In the proposed LAPT framework, two core components are emphasized: knowledge transfer and principle adaptation. To showcase their efficacy, two different verions of LAPT are performed on Trans101. In the first version, the principle adaptation step is omitted, referred to as WO Adaptation. In this scenario, the search space is refined solely based on the design principles derived from previous tasks and never refined during the search process. In the second version, denoted as WO Transfer, architectural knowledge from prior tasks is unavailable, and LAPT refines the search space solely based on the architectures designed for the new task. The model ranks of the discovered architectures via these two versions illustrated in Fig. 4. These results demonstrate that knowledge transfer closely aligns with enhanced search performance. Without employing the transfer strategy, identifying top-ranking architectures becomes challenging. Furthermore, the principle adaptation strategy is more related to convergence of the search method (leading to the lower variance of model ranks).

### v2

@@ Abstract
Transferable neural architecture search (TNAS) has been introduced to design efficient neural architectures for multiple tasks, to enhance the practical applicability of NAS in real-world scenarios. In TNAS, architectural knowledge accumulated in previous search processes is reused to warm up the architecture search for new tasks. However, existing TNAS methods still search in an extensive search space, necessitating the evaluation of numerous architectures. To overcome this challenge, this work proposes a novel transfer paradigm, i.e., design principle transfer. In this work, the linguistic description of various structural components’ effects on architectural performance is termed design principles. They are learned from established architectures and then can be reused to reduce the search space by discarding unpromising architectures. Searching in the refined search space can boost both the search performance and efficiency for new NAS tasks. To this end, a large language model (LLM)-assisted design principle transfer (LAPT) framework is devised. In LAPT, LLM is applied to automatically reason the design principles from a set of given architectures, and then a principle adaptation method is applied to refine these principles progressively based on the new search results. Experimental results show that LAPT can beat the state-of-the-art TNAS methods on most tasks and achieve comparable performance on others. Code is available at https://github.com/milkmilk511/LAPT.
@@ Introduction
  • An LLM-assisted framework is proposed to implement the design principle transfer across different NAS tasks, which offers at least three advantages: (i) Learning of the general design principles based on LLMs; (ii) Task-specific principle adaptation against domain shit; (iii) Improved interpretability of search space refinement.
@@
To reduce the computational cost, transfer learning has been incorporated into NAS, giving rise to TNAS (Huang et al. 2022; Lee and Hyung 2021; Lu et al. 2020, 2021; Elsken et al. 2020; Zhou et al. 2023). In TNAS, the knowledge gained during the architecture search for one task is applied to aid in the architecture design for others. Depending on the types of transferred knowledge, TNAS can be classified into two categories. One category transfers already found architectures from previous tasks to directly build architectures for new tasks. For instance, Lu et al. (Lu et al. 2020) design architectures for CIFAR-100 and use them to classify images from ImageNet directly. However, the generalization performance of these methods has been criticized. To solve this drawback, multitasking evolutionary NAS (MTNAS) (Zhou et al. 2023) applies these high-performing architectures to guide the architecture search for new tasks instead of solving them directly.
The second category reuses the NAS model built in previous tasks to guide the architecture search for a new task. For example, Lu et al. (2021) leverage the supernet that has been initialized on ImageNet to improve the search efficiency of evolutionary NAS (ENAS) on new image classification tasks. In another related work (Elsken et al. 2020), a meta-architecture is acquired from previous tasks, and this architecture can be adapted to a new task through just a few gradient-based steps. Huang et al. (2022) fine-tune the performance predictor developed for the Jigsaw task and use it to reduce the search cost on other computer vision tasks. However, these methods still search in a very large search space, consuming a lot of computing resources for obtaining high-quality architectures. In this work, a new paradigm of TNAS is proposed, i.e., design principle transfer. In this method, general design principles are summarized from established architectures and further reused to reduce the search space for new NAS tasks.
@@
An overview of LAPT is presented in Figure 1 and Algorithm 1. This framework consists of two stages, i.e., design principle learning and the architecture search for new tasks. In the design principle learning stage, a set of well-performing architectures $\mathcal{A}$ is collected from a pre-defined search space $\Omega$ (line 1 in Algorithm 1). Then, a prompt is devised to help a pre-trained LLM reason general design principles $P$ from $\mathcal{A}$ (line 2 in Algorithm 1). These learned principles $P$ are transferred as the initial design principles $\{P_{n}\}_{n=1}^{N}$ to help solve $N$ new NAS tasks $\{\mathcal{T}_{n}\}_{n=1}^{N}$ (lines 3-10 in Algorithm 1). Specifically, for the task $\mathcal{T}_{n}$, $P_{n}$ is translated to the subsets $\{\mathcal{C}^{r}_{i},~{}\mathcal{S}^{r}_{i}\}_{i}^{L}$ by the pre-trained LLM, leading to the generation of a refined search space $\Omega_{r}$. Then, a NAS method is applied to search for the promising architectures $\mathcal{B}_{n}$ from $\Omega_{r}$ for $\mathcal{T}_{n}$. These found architectures are applied to adapt $P_{n}$ to $\mathcal{T}_{n}$. These steps related to principle adaptation and architecture search are repeated until reaching a predefined stopping criterion. Finally, the best architecture found in the search process is used to solve $\mathcal{T}_{n}$.
@@
1:  Build an archive $\mathcal{A}$ consisting of performing architectures from $Omega$;
2:  $\{P_{n}\}_{n=1}^{N}\leftarrow$ Prompt the LLM to learn design principles $P$ from $\mathcal{A}$ and set it as initial principles for each task;
3:  for $n$ from 1 to $N$ do
4:     $Base\leftarrow 0$;
@@
9:        $P_{n},~{}Base\leftarrow$ Adapt $P_{n}$ based on $\mathcal{B}_{n}$, $Base$, $Best$ and return the updated principles; #Alg. 2
@@
Firstly, the pre-trained LLM benefits from exposure to a wide array of programming languages, allowing it to gain awareness of the neural architecture from source codes (Zheng et al. 2023). Nevertheless, due to the token limitation, it becomes infeasible to feed all architecture source codes directly into the LLM. To tackle this issue, Python classes that can instantiate an architecture from its architectural parameters, i.e., ${\cup}_{i=1}^{L}\{l_{i},~{}s_{i}\}$, are set as prompts. This approach enables LLMs to assimilate knowledge about these neural architectures solely through a few architecture parameters. A simple example of the Python code-based prompt is shown in Figure. 2.
@@
In the architecture search stage, principles $P$ learned in the last stage are used to help build the refined search space $\Omega_{r}$. Specifically, we prompt the pre-trained LLM to translate $P$ into the most available candidate operators $\mathcal{C}_{i}$ and information sources $\mathcal{S}_{i}$ for each layer. In this way, $\Omega_{r}$ can be built following equation (3). Then, the NAS method is used to search architectures $\mathcal{B}_{n}$ from $\Omega_{r}$, aiming to solve the problem shown in equation (1).
However, due to domain shift where the architecture performs differently on different tasks, $\Omega_{r}$ may not be optimal for all the tasks. To alleviate the potential negative effects of domain shift, this work adapts $P$ to the target task based on the newly found architectures. As a result, a task-specific subspace can be built progressively. Specifically, the NAS method is applied to search for architecture from $\Omega_{r}$ in an iterative way. In each iteration, if the better-performing architectures are found, LLM is prompted to update $P_{n}$ based on these architectures; otherwise, LLM is required to describe the effects of other available candidate operators and information sources that are not in $P_{n}$, promoting the exploration for other promising regions in $\Omega$. Details of this adaption strategy can be seen in Algorithm 2.
@@
Input: design principles $P_{n}$, found architectures $\mathcal{B}_{n}$, $Base$, $Best$, # of selection architectures $r$.
@@
Trans101. Architectures in Trans101 follow the same cell-based structure in NAS201, but there are only 4 candidate operators for each layer. Thus, only 4K candidate architectures are included in the space.
@@
Experiments on NAS201. In these experiments, we collect 50 top-performing architectures on ImageNet from NAS201 as the input for LLM. Then, learned principles are used to constrain the range of available operators on each layer, leading to a refined search space of NAS201. We separately search for top-performing architectures (i.e., Top 0.1%) on CIFAR-10 and CIFAR-100 from the refined search space and record the time cost.
@@
In these experiments, GPT-4 is used as the pre-trained LLM for design principle learning and adaptation. Related prompts can be seen in Appendix II. In addition, the effectiveness of different LLMs is shown in Appendix V.
Table 2: Comparison with TNAS and NAS methods on NAS201.We present the test accuracy achieved by the found architectures on two unseen datasets. Additionally, we provide the number of neural architectures (Archs.) that are trained in the search. We run LAPT 20 times with different random seeds and average values are reported.
@@
Comparison on NAS201. We choose well-known NAS methods and TNAS methods as baselines: (1) the vanilla NAS method REA and the state-of-the-art NAS method HEBO (Cowen-Rivers et al. 2022); (2) the state-of-the-art TNAS methods include TNAS-BO (Shala et al. 2023) and MetaD2A (Lee and Hyung 2021). In LAPT, REA is used for architecture search, and its results (LAPT-REA) are shown in Table 2.
@@
Compared with TNAS methods, benefiting from the search space refinement, LAPT can achieve the SOTA performance. Specifically, all of these methods can reach the top architecture on CIFAR-10, but LAPT is about 25x faster than MetaD2A and 7x faster than TNAS-BO. On CIFAR-100, LAPT is about 12x faster than MetaD2A obviously, and is about 7x faster than TNAS-BO.
@@
From Table 6, our LAPT can achieve better search performance and search efficiency by a large margin than the NAS methods. Specifically, LAPT can beat BONAS on all the 6 tasks. LAPT outperforms REA on 5 out of 6 tasks and achieves the highest average mode rank.
@@
Table 4: Performance comparison on CIFAR-10 using the search space DARTs.
@@
Search results on ImageNet are shown in Table 5. From this table, we can see that because of the search space refinement, LAPT can achieve similar search results with other NAS methods but with about 1.5-2$\times$ less search cost (2 GPU days), which is more time-efficient. The learned general principles for DARTS are shown in Appendix IV.
@@
In the proposed LAPT framework, two core components are emphasized: knowledge transfer and principle adaptation. To showcase their efficacy, two different verions of LAPT are performed on Trans101. In the first version, the principle adaptation step is omitted, referred to as WO Adaptation. In this scenario, the search space is refined solely based on the design principles derived from previous tasks and never refined during the search process. In the second version, denoted as WO Transfer, architectural knowledge from prior tasks is unavailable, and LAPT refines the search space solely based on the architectures designed for the new task. The model ranks of the discovered architectures via these two versions are illustrated in Fig. 4. These results demonstrate that knowledge transfer closely aligns with enhanced search performance. Without employing the transfer strategy, identifying top-ranking architectures becomes challenging. Furthermore, the principle adaptation strategy is more related to the convergence of the search method (leading to the lower variance of model ranks).
@@ Conclusion
Acknowledgments
This work was supported in part by National Key R&D Program of China (2022YFC3801700); in part by the Research Grants Council of the Hong Kong SAR (Grant No. PolyU11211521, PolyU15218622, PolyU15215623, and C5052-23G), and the National Natural Science Foundation of China (Grant No. U21A20512).
Appendix I Additional Related Work
Neural architecture search
Neural architecture search (NAS) has made significant strides in the automatic construction of deep neural network (DNN) architectures (Zhou et al. 2021b). NAS typically follows an iterative framework: a search method is utilized to discover promising architectures within a predefined search space. Then, through a tailored performance evaluation system, evaluations of these architectures serve as feedback to guide the search toward better-performing architectures.
Three primary NAS methods have been introduced in existing works, i.e., reinforcement learning-based methods (denoted as RLNAS) (Baker et al. 2016; Zoph and Le 2016), gradient-based NAS methods (denoted as GDNAS) (Liu, Simonyan, and Yang 2018), and evolutionary algorithm-based methods (denoted as ENAS) (Sun et al. 2018). RLNAS trains an agent to build an architecture layer by layer following a decision-making process. In ENAS, each individual corresponds to an architecture, and the population consisting of these individuals evolves via genetic operations to discover the optimal architecture. During the search process of both RLNAS and ENAS, massive architectures are required to be trained and evaluated, resulting in expensive time cost. To alleviate this challenge, GDNAS methods reformulate NAS as a continuous optimization problem and employ gradient-based methods to drive the search process. However, due to the reliance on gradient information, GDNAS methods often incur high computational memory demands and local optimal neural architectures.
Figure 5: Prompts used to drive LLMs to learn and adapt the design principles from architectures in NAS201.
Figure 6: Prompts used to drive LLMs to learn and adapt the design principles from architectures in Trans101.
Figure 7: Prompts used to drive LLMs to learn and adapt the design principles from architectures in DARTs.
Design principles of DNNs
Recently, design in a principled way has been proposed to reduce the search space in NAS (Radosavovic et al. 2020; Shen et al. 2023). In these approaches, design principles, which describe how various structural components (such as pooling, convolution, and their interconnections) impact architecture performance, are leveraged to optimize the search space. Specifically, architectures containing less critical components are pruned from the search space according to these principles, resulting in a more refined subspace. This refined subspace contains a higher proportion of high-performing architectures than the original search space (Radosavovic et al. 2020). Exploration within this refined search space significantly enhances the search effectiveness and efficiency. For instance, by incorporating principles like uniform stage depth and a non-decreasing number of channels, a CNN search space can be condensed from 1010 to 102, enabling the discovery of top-performing architectures within seconds (Wang et al. 2024). However, learning design principles heavily relies on expert knowledge. Additionally, given the intricate and diverse nature of DNN architectures, it is challenging for experts to learn general design principles. These problems restrict the practical application of design principles in real-world scenarios.
Appendix II Prompt Engineering
In this work, we design specific prompts to help LLMs reason the general design principles from given architectures, translate these principles into architectural parameters, and adapt these principles to specific tasks. These prompts consist of one or several parts:
  1. 1.
    Task description: It informs LLMs of the problem that is to be solved.
  2. 2.
    Strategy: It informs LLMs of the steps that should be followed to solve the given problem.
  3. 3.
    Expected output: It informs LLMs of the type of information that they should output.
  4. 4.
    Note: It provides additional instructions for LLMs to improve their efficiency.
Prompts used in different spaces share a similar template but with a few differences among their contents. Details of these prompts are shown in Figure 56, and 7.
(a) Obj
(b) SC
(c) Roo
(d) Auto
(e) Nor
(f) Seg
Figure 8: Quality of search spaces on Trans101.
Appendix III Quality Analysis of Search Space
In this work, the proposed framework LAPT optimizes the search space by pruning the unpromising architectures from it. In this way, a refined search space with a higher proportion of well-performing architectures can be built. To demonstrate the effectiveness of this method, we measure the quality of these search spaces from Transfer101 and present these results in this section.
Specifically, the error empirical distribution function (eEDF) is used to measure the quality of the search space, which is shown in equation (1):
$F(e)=\frac{1}{n}\sum_{i=1}^{n}{1[e_{i}\leq e]}.$ (4)
$F(e)$ gives the fraction of models with errors less than $e$, and $n$ is the total number of architectures in the search space or the subspace. We conduct the proposed LAPT on Trans101 to solve the 6 computer vision tasks (i.e., Object Classification (Obj), Scene Classification (SC), Room Layout (Roo), AutoEncoder (Auto), Surface Normal (Nor), Semantic Segmentation (Seg)) based on the architectural knowledge from the prior task Jigsaw. For each task, we measure the quality of the entire search space (termed original), the refined search space based on the general design principles (termed iteration_0), and the task-specific search space built based on the adapted design principles (termed iteration_3). The measurements are shown in Figure 8.
From these results, these task-specific search spaces have the highest proportion of well-performing architectures. Then, the ones refined based on the general design principles are also of higher quality than the original one. These results demonstrate that: (1) the proposed LLM-based strategy can well learn the useful general design principles that can be used to identify the unpromising/promising architectures; (2) the design principle transfer paradigm is effective for the search space refinement in new NAS tasks; (3) our principle adaptation is effective in building the task-specific search space which is more effective for the specific NAS task.
Appendix IV General Design Principles
In this work, we drive the pre-trained LLMs to learn design principles for different search spaces from established architectures. These general design principles can be reused to prune the search space, leading to a refined search space for the new NAS tasks. The design principles of the architecture in NAS201 have been shown in the main paper. In this part, we present the design principles of the search spaces Trans101 and DARTs, respectively.
Figure 9: General design principles on Trans101.
Firstly, the design principles for Trans101 are shown in Figure 9. In this search space, these principles mainly describe the influence of different operators (e.g., pooling layer, skip, and convolution), because the number of layers and connections in these architectures are fixed. These principles first point out convolution layers should be used in the early layers of the architecture to extract feature information, and the skip operator should be used in the middle layers to prevent gradient vanishing issues. This is identical to the expert knowledge on convolution neural networks (CNNs). Additionally, the usage of other layers is also presented.
Figure 10: General design principles on Trans101.
Table 6: Mode rank of the found architecture on each task.
LLM Obj SC Roo Auto Nor Seg Ave.
GPT-3.5 40.3$\pm$8.1 16.7$\pm$3.9 61.0$\pm$8.3 28.3$\pm$5.6 51.2$\pm$10.4 61.9$\pm$11.2 43.2
GPT-4 9.0$\pm$3.0 4.2$\pm$2.5 14.4$\pm$4.1 25.7$\pm$2.8 14.7$\pm$3.5 6.0$\pm$1.8 12.3
GPT-4o 9.9$\pm$2.9 8.8$\pm$4.0 9.9$\pm$2.8 23.3$\pm$3.0 9.3$\pm$2.3 11.4$\pm$3.0 12.1
Secondly, the design principles for DARTs are shown in Figure 10. In this search space, these principles mainly describe the available operators in each stage and its information source. Because two kinds of cells should be designed, i.e., the normal cell and the reduction cell, we present the design principles of them separately. Specifically, for the normal cell, the complexity of the operators increases with the cell’s depth, where the high-complexity operators such as convolution are used to extract feature information in the early stage and the low-complexity operators such as skip and pooling are used in the later stage to prevent gradient vanishing. However, the reduction cell is designed in a different way from the normal one. In the reduction cell’s early stage, the low-complexity operators such as pooling are used for the feature size reduction, and high-complexity operators are used in the later stages for feature extraction. The reason is that these cells work in different mechanisms. As for the information sources, all the stages intend to receive information from the previous one to build a deep structure. More details can are in Figure 10.
Appendix V Effectiveness of LLMs
In our proposed LART framework, the pre-trained LLMs play an important role in design principle learning, design principle adaptation, and design principle translation. This part further discusses the effectiveness of LLMs on the performance of LAPT. To this end, we conduct LLMs with different pre-trained LLMs on Trans101, and the model ranks of architectures found for each task are shown in Table I.
From these results, GPT-4 and GPT-4o achieve comparable performance on these tasks, and both of them outperform GPT-3.5 by a large margin. The reason is that GPT-4 and GPT-4o have more powerful language reasoning, understanding, and generation capabilities. It demonstrates that the performance of the proposed LAPT framework is close to the used LLMs.